OpenStudy (anonymous):

can someone help me write the equation of the line which passes through (2, 1) and is perpendicular to x = –2 ?

OpenStudy (unklerhaukus):

so you want the form \[y(x)=mx+b\]where the slope is \(m\)

OpenStudy (unklerhaukus):

actually you dont want that form

OpenStudy (anonymous):

y-1=m(x-x1) I dont know I have to use this formula because there are problems in wich u use this but the are problem in which you dont use it

OpenStudy (unklerhaukus):

yeah the formula dosent work for vertical lines

OpenStudy (unklerhaukus):

x=-2 is a vertical line that intercepts at (-2,0) on the x axis

OpenStudy (anonymous):

Am actually cofused how do I know when they are vertical or horizontal . So this one is vertical??

OpenStudy (unklerhaukus):